# Reciprocal Beyond-Diagonal Reconfigurable Intelligent Surface  (BD-RIS):  Scattering Matrix Design via Manifold Optimization

### v3

@@ III-A Solution for the Group-Connected architecture
1: Compute initial objective $\eta(\mathbf{\Theta}^{(0)})$ and Riemannian gradient $\mathbf{r}^{(0)}=\mathrm{T}_{\mathbf{\Theta}}(\nabla_{\mathbf{\Theta}}\eta(\mathbf{\Theta}^{(0)}),\mathbf{\Theta}^{(0)})$
2: Set initial search direction $\boldsymbol{\Xi}^{(0)}=-\mathbf{r}^{(0)}$
3:for $i=0$ to II do
4:  if $\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle\leq 0$ then
5:   Set $\boldsymbol{\Xi}^{(i)}=\mathbf{r}^{(i)}$
6:  end if
@@
7:  Compute $\alpha^{(i)}$ via Armijo line search and update $\mathbf{\Theta}^{(i+1)}$
8:  Compute new objective $\eta(\mathbf{\Theta}^{(i+1)})$, using (16a)
9:  Compute new Riemannian gradient $\mathbf{r}^{(i+1)}$, using (39)
@@
10:  Compute $\beta^{(i)}=\max\left(0,\frac{\langle\mathbf{r}^{(i+1)},\mathbf{r}^{(i+1)}-\mathbf{r}^{(i)}\rangle}{\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle}\right)$
11:  Update direction: $\boldsymbol{\Xi}^{(i+1)}=-\mathbf{r}^{(i+1)}+\beta^{(i)}\boldsymbol{\Xi}^{(i)}$
12:  if $|\eta(\mathbf{\Theta}^{(i+1)})-\eta(\mathbf{\Theta}^{(i)})|<\epsilon$ then
13:   break
14:  end if
15:end for
16:for $g=1$ to $G$ do
17:  Extract block $\mathbf{\Theta}_{g}$ from $\mathbf{\Theta}^{(i+1)}$
18:  Symmetrize: $\mathbf{\Theta}_{\mathrm{sym}}=\frac{1}{2}(\mathbf{\Theta}_{g}+\mathbf{\Theta}_{g}^{T})$, using (40)
19:  Perform SVD: $\mathbf{\Theta}_{\mathrm{sym}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}$, using (41)
20:  Set $\mathbf{\Theta}^{\mathrm{opt}}_{g}=\mathbf{U}\mathbf{V}^{H}$, using (42)
21:end for
22: Assemble full $\mathbf{\Theta}^{\mathrm{opt}}$ as block diagonal of all $\mathbf{\Theta}_{g}^{\mathrm{opt}}$
23:return $\mathbf{\Theta}_{\mathrm{opt}}$
@@
The full procedure implemented in this article is outlined in Algorithm 1. The algorithm begins by initializing a random unitary symmetric block-diagonal matrix $\mathbf{\Theta}^{(0)}$ and computing the initial objective function along with its gradient. In each iteration, these quantities are updated according to the current values of the scattering matrix until convergence is achieved. The step size $\alpha$ is determined using an Armijo line search [40, 41], which ensures that there is a sufficient ascent. Furthermore, the parameter $\beta$ denotes the conjugate gradient (CG) momentum term, used to enhance the convergence speed of the algorithm.
@@
Among various options, including the Fletcher-Reeves, Polak Ribière, and Hestenes-Stiefel formula, this algorithm incorporates the Polak-Ribière formula as described in [37, 41]. Finally, the scattering matrix resulting from the series of retraction and tangential projection operations is further processed using the method described in (40)-(42), ensuring that $\mathbf{\Theta}_{g}^{\mathrm{opt}}$ satisfies both the symmetric and unitary constraints.
@@
Computational Complexity: The computational complexity of Algorithm 1 is primarily influenced by the Armijo line search and gradient computation. Within the line search, two operations dominate the complexity, namely a triple matrix multiplication and a QR decomposition-based retraction. Considering the general case of the group-connected architecture, the QR decomposition for each group, of size $R/G$, has a complexity of $\mathcal{O}((R/G)^{3})$, while the triple matrix multiplication has a complexity of $\mathcal{O}(KR(R+N))$. Since the QR decomposition is applied independently to each group, the total complexity of the retraction operation across all groups and users is given as $\mathcal{O}(KG(R^{3}/G^{3}))$.
@@
In contrast, the triple matrix multiplication is performed on the full system, $i.e.$, not on a per-group basis, and yields a total complexity for all users of $\mathcal{O}(K(KR^{2}+KRN))$. With $L$ Armijo iterations, these contributions accumulate, giving an overall line search complexity of
$\mathcal{O}(LKG(R^{3}/G^{3})+LK(KR^{2}+KRN)).$ (43)
The gradient computation consists of several terms, resulting in a total complexity across all groups and users of
$\hskip-8.61108pt\mathcal{O}(KG(N^{2}\!+\!N\!R/G\!+\!(R/G)^{2}\!+\!K\!(R/G)^{2})\!+\!G(R/G)^{2}).\vskip-2.15277pt$ (44)
@@
Furthermore, the initialization and projection described in (40)-(42) have a computational complexity of $\mathcal{O}(R^{3}/G^{2})$, which stems from the SVD computation. Lastly, the tangential projection in (39) has a complexity of $\mathcal{O}(R^{3})$. Considering $I$ iterations of the algorithm, and the case $L\gg 1$, the total complexity of the algorithm can then be expressed as
$\mathcal{O}(IL(KR^{3}/G^{2}+K^{2}R^{2}+K^{2}RN)).\vskip-2.15277pt$ (45)
@@
Table I summarizes the simplified computational complexity associated with the different BD-RIS architectures, allowing for a straightforward comparison with SotA methods [^5] Notice that the SotA method in [26] assumes that the number of TX antennas at the BS equals the number of single-antenna users ($K=N$). Our method, does not require such assumption and therefore can be considered a generalization of the latter in this regard., where $t$ denotes the number of iterations required by the SotA algorithm [26]. The convergence behavior of Algorithm 1 is shown in Figure 2 for the different BD-RIS architectures considered, $i.e.$, single-, group-, with group size of 2 and 4, and fully-connected, respectively.
@@
Figure 2: Convergence of Algorithm 1 vs. number of iterations II for different connectivity structures, with $P_{\mathrm{max}}=20\,\mathrm{dBm}$, $K=5$, $N=5$, and $R=32$.
@@
For clarity, each BD-RIS architecture is represented by a distinct marker, as indicated in the legend of Figure 2, and this notation is used consistently throughout the article to facilitate cross-referencing of results.
@@
The simulation settings are as follows: convergence tolerance of $\epsilon=10^{-8}$, maximum number of CGA iterations $I=8000$, maximum Armijo line search steps $L=200$, sufficient increase coefficient for the stepsize of $2\times 10^{-11}$, initial stepsize $c_{\text{init}}=1$, contraction factor for the stepsize $c_{\text{dec}}=0.75$, and the symmetry enforcing penalty constant $\nu=1$. Results indicate that more complex architectures require a larger number of iterations to converge; for instance, the single-connected case converges in approximately $50$ iterations, while the full-connected case requires around $2500$ iterations under the same conditions. Additionally, the convergence speed of the algorithm depends on the number of REs and the maximum transmit power $P_{\mathrm{max}}$, as well as on the initial scattering matrix, which in this article is chosen as a random unitary symmetric matrix. The figure represents results generated for one Monte-Carlo iteration, hence the reason for the non-smooth convergence curves.
@@
As validated in Figure 2, the number $I$ of iterations required for convergence is considered for the scenario where $P_{\mathrm{max}}=20\mathrm{dBm}$, $K=5$ users, $N=5$ TX antennas, and $R=32$ REs. Specifically, these are $I_{\mathrm{SC}}=50$, $I_{\mathrm{GC}_{2}}=650$, $I_{\mathrm{GC}_{4}}=1050$ and $I_{\mathrm{FC}}=2500$, corresponding to the single-, group-, with group size 2 and 4, and fully-connected architectures, respectively.
@@
As can be learned from Table I, our proposed algorithm offers scalability, since its computational complexity grows with lower exponents on key system parameters such as the number of users $K$, the number of RIS elements $R$ and the number of BS antennas $N$. This advantage is in addition to the superior communications performance, which shall be throughly demonstrated in the next section. For systems of smaller sizes, however, the performance advantage of the proposed method may come at the cost of a higher complexity compared to the SotA technique [26], due to the large number of iterations required for convergence of the algorithm, as shown in Figure 2. The improved sum-rate and increased robustness observed across multiple BD-RIS architectures, resulting from the algorithm’s design, are thoroughly presented and analyzed in the subsequent section.
@@ IV Simulation Results
This section aims to showcase the effectiveness of the proposed BD-RIS-aided system in enhancing communication performance via computer simulations executed with the following parameters. I will be assumed throughout the experiments that the BS is equipped with $N=5$ TX antennas, serving $K=5$ single RX antenna users, the RIS structures consist of $R=32$ REs and the maximum transmit power $P_{\mathrm{max}}$ ranges from 0 to $20\,\mathrm{dBm}$. Furthermore, the parameters utilized in Algorithm 1 are the same as those previously used to generate the results shown in Figure 2. In addition to the performance analysis of the scattering matrix design presented in this article, the results also showcase the sum-rate behavior of our proposed scattering matrix in combination with different SotA beamforming methods [42].
In accordance with [26], the channel model adopted in this article assumes i.i.d. Rayleigh fading with distance-dependent large-scale fading expressed as $\Upsilon(d)=C_{0}(\frac{d}{d_{0}})^{-\rho}$, where $C_{0}=-30\mathrm{dB}$ is the reference channel gain at a reference distance $d_{0}=1$m, $\rho=2.2$ is the path-loss exponent, and d=50d=50m is the distance between the BS and BD-RIS  [^6] Although i.i.d fading channels are, strictly speaking, inconsistent with two-dimensional (2D) surfaces, as shown in [43], we emphasize that this particular channel model is adopted here merely in order to facilitate a fair comparison with the SotA method [26], since the scattering matrix design itself does not require any specific model to be feasible.. The system is assumed to operate at 2.42.4GHz, and each user experiences identical noise with a noise power $N_{0}=-80\mathrm{dBm}$.
The first set of results is presented in Figure 3 to demonstrate the effectiveness of the proposed method in enhancing system performance. Different architectures, such as the single-, group-, and fully-connected are considered, and results comparing the sum-rate performance achieved by the proposed and the SotA scattering matrix design [26] are provided. Notably, the results indicate that the proposed scattering matrix outperforms the SotA under the same beamforming conditions, $i.e.$, using uniform power allocation. In particular, as seen in Figure 3a, the proposed single-connected scattering matrix design surpasses both the SotA single- and group-connected $(R_{G}=2)$ architectures.
@@
Figure 3: Comparison of sum-rate performance of the proposed vs. SotA [26] scattering matrix design with uniform power allocation, considering the single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
Figure 4: CDF of sum-rate performance of the proposed vs. SotA [26] scattering matrix design with uniform power allocation, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
Complementing the assessment in terms of sum-rate performance, Figure 4 further shows the cumulative distribution function (CDF) of the sum-rate performance of the proposed BD-RIS scattering matrix design, alongside the SotA method [26] for all considered architectures, $i.e.$, single-, group- with group sizes of 2 and 4, and fully-connected.
The results show that, for architectures of comparable complexity, the proposed scattering matrix design consistently outperforms the SotA design. Additionally, it is observed that the single-connected architecture using the proposed method achieves performance comparable to the group-connected architecture with a groups size of 2, which has an order of magnitude higher complexity in this scenario.
Next, Figure 5 presents a comparison of the sum-rate performance of the proposed scattering matrix with different beamforming schemes, namely uniform power allocation “UP”, FP-based beamforming “BF” and power allocation “PA” [42], against the SotA scattering matrix with “UP” [26]. It is observed that the use of the FP-based TX beamformer further improves the sum-rate performance. In particular, the sum-rate achieved by the proposed single-connected architecture in combination with the SotA FP-based beamforming “BF” matrix [42] is emphasized.
Finally, Figure 6 presents the sum-rate performance of the proposed method in comparison to the SotA scattering matrix design technique of [26], in combination with uniform power allocation “UP”, as a function of the number of REs $R$. This result further builds upon the claim that for all BD-RIS architectures considered, $i.e.$, single-, group-, and fully-connected, and under every evaluated scenario the proposed scattering matrix consistently outperforms the SotA benchmark [26]. Importantly, the performance gain is maintained as the number of reflective elements increases, which highlights the robustness of the proposed approach, and showcases that the proposed solution is scalable.
@@
Figure 5: Comparison of sum-rate performance using different beamforming schemes, considering the proposed and SotA [26] single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
Figure 6: Comparison of sum-rate performance of the proposed vs. SotA [26] scattering matrix designs as a function of the number of REs, i.e. $R$, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.

### v4

@@ III-A Solution for the Group-Connected Architecture
2: Compute initial objective ${\breve{\eta}}(\mathbf{\Theta}^{(0)})$ and Riemannian gradient $\mathbf{r}^{(0)}=\mathrm{T}_{\mathbf{\Theta}}(\nabla_{\mathbf{\Theta}}{\breve{\eta}}(\mathbf{\Theta}^{(0)}),\mathbf{\Theta}^{(0)})$ using (42)
3: Set initial search direction $\boldsymbol{\Xi}^{(0)}=-\mathbf{r}^{(0)}$
4:for $i=0$ to II do
5:  if $\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle\leq 0$ then
6:   Set $\boldsymbol{\Xi}^{(i)}=\mathbf{r}^{(i)}$
7:  end if
@@
8:  Compute $\alpha^{(i)}$ via Armijo line search and update $\mathbf{\Theta}^{(i+1)}$
9:  Compute new FP auxiliary variables $\tau_{k}$ and $y_{k}$, $\forall k$
10:  Compute new objective ${\breve{\eta}}(\mathbf{\Theta}^{(i+1)})$, using (25a)
11:  Compute new Riemannian gradient $\mathbf{r}^{(i+1)}$, using (43)
@@
12:  Compute $\beta^{(i)}=\max\left(0,\frac{\langle\mathbf{r}^{(i+1)},\mathbf{r}^{(i+1)}-\mathbf{r}^{(i)}\rangle}{\langle\mathbf{r}^{(i)},\boldsymbol{\Xi}^{(i)}\rangle}\right)$
13:  Update direction: $\boldsymbol{\Xi}^{(i+1)}=-\mathbf{r}^{(i+1)}+\beta^{(i)}\boldsymbol{\Xi}^{(i)}$
14:  if $|\eta(\mathbf{\Theta}^{(i+1)})-\eta(\mathbf{\Theta}^{(i)})|<\epsilon$ then
15:   break
16:  end if
17:end for
18:for $g=1$ to $G$ do
19:  Extract block $\mathbf{\Theta}_{g}$ from $\mathbf{\Theta}^{(i+1)}$
20:  Symmetrize: $\mathbf{\Theta}_{\mathrm{sym}}=\frac{1}{2}(\mathbf{\Theta}_{g}+\mathbf{\Theta}_{g}^{T})$
21:  Perform SVD: $\mathbf{\Theta}_{\mathrm{sym}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H}$
22:  Set $\mathbf{\Theta}^{\mathrm{opt}}_{g}=\mathbf{U}\mathbf{V}^{H}$
23:end for
24: Assemble full $\mathbf{\Theta}^{\mathrm{opt}}$ as block diagonal of all $\mathbf{\Theta}_{g}^{\mathrm{opt}}$
25:return $\mathbf{\Theta}_{\mathrm{opt}}$
@@
The algorithm begins by initializing a random unitary symmetric block-diagonal matrix $\mathbf{\Theta}^{(0)}$ and computing the initial objective function along with its gradient. In each iteration, these quantities are updated according to the current values of the scattering matrix until convergence is achieved. The step size $\alpha$ is determined using an Armijo line search [30, 50], which ensures that there is a sufficient ascent. Furthermore, the parameter $\beta$ denotes the conjugate gradient (CG) momentum term, used to enhance the convergence speed of the algorithm.
@@
Among various options, including the Fletcher-Reeves, Polak Ribière, and Hestenes-Stiefel formula, this algorithm incorporates the Polak-Ribière formula as described in [47, 50]. Finally, the scattering matrix resulting from the series of retraction and tangential projection operations is further processed using the method described in [22], ensuring that $\mathbf{\Theta}_{g}^{\mathrm{opt}}$ satisfies both the symmetric and unitary constraints.
@@
Computational Complexity: The computational complexity of Algorithm 1 is primarily influenced by the Armijo line search. Within the line search, two operations dominate the complexity, namely a triple matrix multiplication and a QR decomposition-based retraction. Considering the general case of the group-connected architecture, the QR decomposition for each group, of size $R/G$, has a complexity of $\mathcal{O}\big((R/G)^{3}\big)$, while the triple matrix multiplication has a complexity of $\mathcal{O}\big(KR(R+N)\big)$. Since the QR decomposition is applied independently to each group, the total complexity of the retraction operation across all groups and users is given as $\mathcal{O}\big(KG(R^{3}/G^{3})\big)$.
@@
In contrast, the triple matrix multiplication is performed on the full system, $i.e.$, not on a per-group basis, and yields a total complexity for all users of $\mathcal{O}\big(K(KR^{2}+KRN)\big)$. With $L$ Armijo iterations, these contributions accumulate, giving an overall line search complexity of
$\mathcal{O}\big(LKG(R^{3}/G^{3})+LK(KR^{2}+KRN)\big).\vskip-2.15277pt$ (44)
@@
The gradient computation consists of several terms, resulting in a total complexity across all groups and users of $\mathcal{O}\big(K^{2}G(R/G)^{2}\big)$, and the computation of the FP auxiliary variables, which further amounts to $\mathcal{O}\big(K^{2}N\big)$. Furthermore, the initialization and projection described in [22] have computational complexity $\mathcal{O}\big(R^{3}/G^{2}\big)$, which stems from the singular value decomposition (SVD) computation. Lastly, the tangential projection in (43) has a complexity of $\mathcal{O}\big(R^{3}\big)$. Considering $I$ iterations of the algorithm, and the case $L\gg 1$, the total complexity of the algorithm can then be expressed as
$\mathcal{O}\big(IL(KR^{3}/G^{2}+K^{2}R^{2}+K^{2}RN)\big).\vskip-2.15277pt$ (45)
@@
Table I summarizes the simplified computational complexity associated with the different BD-RIS architectures, enabling a direct comparison with the SotA method from [21] [^12] Notice that the aforementioned SotA technique assumes that the number of TX antennas at the BS equals the number of single-antenna users ($K=N$). The proposed method does not require such assumption and therefore can be considered a generalization of the latter in this regard., where $t$ denotes the number of iterations required by the interference nulling algorithm [21]. In addition, we compare the proposed method against the joint scattering and beamforming matrix design approach based on the pp-ADMM framework [34]. Accounting for $T_{\mathrm{ADMM}}$ iterations required to satisfy the convergence criterion, the overall computational complexity of the pp-ADMM algorithm scales as
$\mathcal{O}\Big(T_{\mathrm{ADMM}}\Big[R^{3}K+\big(\min\!\big\{\sum\nolimits_{i=1}^{R}|S_{i}|,\,2RK\big\}\big)^{3}\Big]\Big),\vskip-2.15277pt$ (46)
where $S_{i}$ denotes the set of non-zero entries ($i.e.$, active connections) in the $i$-th row of the BD-RIS scattering matrix, such that $\sum_{i=1}^{R}|S_{i}|$ represents the total number of independent optimization variables determined by the connectivity structure of the BD-RIS.
@@
Figure 3: Convergence of Algorithm 1 vs. number of iterations II for different connectivity structures, with $P_{\mathrm{max}}=20\,\mathrm{dBm}$, $K=2$, $N=2$, $R=32$, and $K$ factor $K_{f}=$-$\infty$ dB.
The convergence behavior of Algorithm 1 is shown in Figure 3 for the different BD-RIS architectures considered, $i.e.$, single-, group-, with group size of 2 and 4, and fully-connected, respectively. For clarity, the figure represents results generated for one Monte-Carlo iteration, hence the reason for the non-smooth convergence curves.
@@
Each BD-RIS architecture is represented by a distinct marker, as indicated in the legend of Figure 3, and this notation is used consistently throughout the article to facilitate cross-referencing of results.
@@
The simulation settings are as follows: convergence tolerance of $\epsilon=10^{-8}$, maximum number of CGA iterations $I=8000$, maximum Armijo line search steps $L=200$, sufficient increase coefficient for the stepsize of $2\times 10^{-11}$, initial stepsize $c_{\text{init}}=1$, contraction factor for the stepsize $c_{\text{dec}}=0.75$, and the symmetry enforcing penalty constant $\nu=1$. Results indicate that more complex architectures require a larger number of iterations to converge; for instance, the single-connected case converges in approximately $100$ iterations, while the full-connected case requires around $600$ iterations under the same conditions.
Additionally, the convergence speed of the algorithm depends on the number of REs and the maximum transmit power $P_{\mathrm{max}}$, as well as on the initial scattering matrix, which in this article is chosen as a random unitary symmetric matrix. Regarding solution quality, since the reformulated optimization problem in (25a) is convex, the FP-based approach applied to the originally non-convex problem is guaranteed to converge to a stationary point, as the adopted approximation yields a monotonically non-decreasing objective sequence [45, 46]. However, due to the non-convexity of the original problem, the stationary point generally corresponds to a local optimum.
@@
As validated in Figure 3, the number $I$ of iterations required for convergence is considered for the scenario where $P_{\mathrm{max}}=20\mathrm{dBm}$, $K=2$ users, $N=2$ TX antennas, and $R=32$ REs, under Rayleigh fading conditions, $i.e.$, with a $K$-factor $(K_{f})$ corresponding to -$\infty$ dB. Specifically, these are $I_{\mathrm{SC}}=100$, $I_{\mathrm{GC}_{2}}=200$, $I_{\mathrm{GC}_{4}}=400$ and $I_{\mathrm{FC}}=600$, corresponding to the single-, group-, with group size 2 and 4, and fully-connected architectures, respectively.
@@
As can be learned from Table I and (46), our proposed algorithm offers scalability, since its computational complexity grows with lower exponents on key system parameters such as the number of users $K$, the number of RIS elements $R$ and the number of BS antennas $N$. This advantage is in addition to the superior communications performance, which shall be thoroughly demonstrated in the next section.
For systems of smaller sizes, however, the performance advantage of the proposed method may come at the cost of a higher complexity compared to the SotA techniques [21, 34], due to the large number of iterations required for convergence of the algorithm, as shown in Figure 3. The improved sum-rate and increased robustness observed across multiple BD-RIS architectures, resulting from the algorithm’s design, are thoroughly presented and analyzed in the subsequent section.
@@ IV Simulation Results
This section aims to showcase the effectiveness of the proposed BD-RIS-aided system in enhancing communication performance via computer simulations executed with the following parameters. Unless otherwise specified, it will be assumed throughout the simulations that the BS is equipped with $N=2$ TX antennas and serves $K=2$ single RX antenna users [^13] This choice is motivated by the goal of isolating and evaluating the performance gains introduced by the BD-RIS. In more heavily loaded scenarios, these gains may be partially masked by additional beamforming gains., the RIS structures consist of $R=32$ REs and the maximum transmit power $P_{\mathrm{max}}$ ranges from 0 to $20\,\mathrm{dBm}$. Furthermore, the parameters utilized in Algorithm 1 are the same as those previously used to generate the results shown in Figure 3. In addition to evaluating the proposed scattering matrix design, the numerical results present the achieved sum-rate when the optimized BD-RIS is combined with a fixed MMSE-based beamformer. Joint optimization of the scattering matrix and beamforming is not considered, allowing the impact of the proposed scattering matrix design to be assessed independently.
@@
Figure 4: Comparison of sum-rate performance of the proposed vs. SotA [21, 34] scattering matrix design with uniform power allocation, considering the single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
The channel model adopted in this article considers both i.i.d. Rayleigh and Rician small-scale fading with distance-dependent large-scale attenuation expressed as $\Upsilon(d)=C_{0}(\frac{d}{d_{0}})^{-\rho}$, where $C_{0}=-30\,\mathrm{dB}$ denotes the reference channel gain at a reference distance $d_{0}=1$ m, $\rho=2.2$ is the path-loss exponent, $d=50$ m is the distance between the BS and BD-RIS, and $d=2.5$ m is the distance between the BD-RIS and the users [^14] Although i.i.d fading channels are not entirely consistent with two-dimensional (2D) surfaces, as shown in [51], this model is adopted here only to enable a fair comparison with the SotA methods [21, 34], since the scattering matrix design itself does not require any specific model to be feasible. The fading parameters and large-scale attenuation settings are selected consistently with the aforementioned SotA methods to ensure meaningful performance comparisons. Importantly, the proposed scattering matrix design framework does not rely on any specific small-scale fading assumption.. Specifically, some numerical results assume Rician fading with $K_{f}=2\,\mathrm{dB}$, while others consider the Rayleigh case corresponding to $K_{f}=-\infty\,\mathrm{dB}$. Lastly, the system is assumed to operate at 2.42.4GHz, and each user experiences identical noise with a noise power $N_{0}=-80\mathrm{dBm}$.
The first set of results is presented in Figure 4 to demonstrate the effectiveness of the proposed method in enhancing system performance. Different architectures, such as the single-, group-, and fully-connected are considered, and results comparing the sum-rate performance achieved by the proposed and the SotAtechniques [21, 34] are provided. Notably, the results indicate that the proposed scattering matrix “SM”, combined with either uniform power allocation “PA”, or a fixed MMSE-based beamformer “BF”, outperforms both SotAcounterparts considered.
In particular, the interference nulling SM design with uniform PA [21] and the joint scattering and beamforming matrix design based on the pp-ADMM framework [34], are outperformed with all group architectures, under Rayleigh fading conditions with $K_{f}=$-$\infty$ dB and for a small number of users and TX antennas at the BS, i.e., $K=2$ and $N=2$.
Figure 5: CDF of sum-rate performance of the proposed vs. SotA [21, 34] scattering matrix design with uniform power allocation, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
@@
Figure 6: Comparison of sum-rate performance using different beamforming schemes, considering the proposed and SotA [21, 34] single-connected “SC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the fully-connected “FC” architecture.
Complementing the assessment in terms of sum-rate performance, Figure 5 further shows the cumulative distribution function (CDF) of the sum-rate performance of the proposed BD-RIS scattering matrix design, alongside the SotA methods [21, 34] for all considered architectures, $i.e.$, single-, group- with group sizes of 2 and 4, and fully-connected.
More specifically, the group-connected architecture with group size 4, when using the proposed method, clearly outperforms the considered SotA designs with the same architecture. The fully-connected architecture with the proposed method achieves, on average, slightly better performance than [34]. Architectures with lower complexity, namely the single-connected architecture and the group-connected architecture with group size 2, achieve slightly lower performance than [34]; however, all proposed architectures still outperform the SotA interference nulling design in [21].
Next, Figure 6 presents a comparison of the sum-rate performance of the proposed scattering matrix “SM”, in combination with either uniform power allocation “PA”, or a fixed MMSE-based beamformer “BF”, against the corresponding SotA schemes, namely the interference nulling SM design with uniform PA [21] and the joint scattering and beamforming matrix design based on the pp-ADMM framework [34].
The comparison is conducted across all considered architectures under Ricean fading conditions with $K_{f}=2$ dB, and for a small number of users and TX antennas at the BS, specifically $K=2$ and $N=2$, highlighting the robustness of the proposed method across different fading conditions. We remark that notable performance gains are particularly observed for the group-connected architecture, with group size 2, where the proposed method outperforms both SotA methods, even those with a higher connectivity, by an order of 2, in the high transmit power regime.
Figure 7: Comparison of sum-rate performance of the proposed vs. SotA [21, 34] methods as a function of the number of REs, i.e. $R$, considering the fully-connected “FC”, group-connected “GC”, with group sizes of 2 and 4, “GC(2)” and “GC(4)”, and the single-connected ”SC” architecture.
Finally, Figure 7 presents the sum-rate performance of the proposed method, in combination with uniform power allocation “PA”, in comparison to the SotA methods [21, 34], as a function of the number of REs $R$. This result further builds upon the claim that for all BD-RIS architectures considered, $i.e.$, single-, group-, and fully-connected, and under the Rayleigh fading scenario, the proposed scattering matrix consistently outperforms the SotA benchmarks [21, 34]. Importantly, the performance gain is maintained as the number of reflective elements increases, which highlights the robustness of the proposed approach, and showcases that the proposed solution is scalable.